He reduced the price 15% to $6800. What was the original price?
Answer by jojo14344(1513)

Well,when you reduced $8000.00 by 15% is equal to $6800.00.
To show,
$8000*0.85=$6800 (you can also $8000-15%=$6800)
Or to get $8000,vice versa
$6800/0.85 = $8000 ---- orig price
thank you,
Jojo